The fraction of students that are interested in a new brand of sneakers is twice the fraction of those that are not. The fraction of students who are not interested is 1/x. Those who expressed no opinion in the sneakers were 10% of the students who were intersted. What percent of students were interested?
Answer by stanbon(75887)

The fraction of students that are interested in a new brand of sneakers is twice the fraction of those that are not. The fraction of students who are not interested is 1/x. Those who expressed no opinion in the sneakers were 10% of the students who were interested. What percent of students were interested?
--------------------------
No interested = 1/x
Interested = 2/x
no opinion = 0.10(2/x)
----------------------------
Equation:
Total number of students = (1/x) + (2/x) + 0.10(2/x)
Total number = (1 + 2 + 0.2)/x = 3.2/x
------------------------------------------
Percent of interested students = (2/x)/(3.2/x) = 2/(3.2) = 62.5%
=====================================================
Cheers,
Stan H.